Graph hyperbola: y squared / 169 - x squared / 16 = 1
Answer by lwsshak3(6761)

Graph hyperbola: y squared / 169 - x squared / 16 = 1
y^2/169-x^2/16=1
This is a hyperbola with vertical transverse axis with center at (0,0):
see graph below:
y=±(169+169x^2/16)^.5

slopes of asymptotes: a/b=13/4
equation of asymptotes:
y=13x/4
and
-13x/4